1. Block's sells 500 bottles of perfume a month when the price is $7. A huge increase in resource costs causes price to rise to $9 and Block's only manages to sell 460 bottles of perfume. The price elasticity of demand is:
1
Expert's answer
2021-05-10T15:27:12-0400

The change in quantity is:"=\frac {(500-460)} {(500+460)}=0.0417"


The change in price;

"=\frac {(7 - 9)}{(7+9)}=0.125"


Elasticity of Demand then ;"=\frac {0.0417}{0.125}=0.33"